darren is 8 years older then his sister. in three years he will be twice as old as she will be then. how old is each now
Answer by thejackal(72)

let his sister's age now be S
such that Darren's age now can be given as D = S + 8
thus
in 3 years D + 3
he will be twice as old as she will be then (S + 3)*2
so
D + 3 = (S + 3)*2
(S + 8) + 3 = 2S + 6
S + 11 = 2S + 6
11 - 6 = S = 5;
D = 5 + 8 = 13